OpenStudy (devonhoward15):

The function f(x) varies inversely with x and f(x)=-10 when x=5. What is the inverse variation equation?

OpenStudy (devonhoward15):

@Daniellelovee @ijlal

OpenStudy (daniellelovee):

y = k/x

OpenStudy (ijlal):



OpenStudy (daniellelovee):

-10=k/5

OpenStudy (ijlal):

Whenever the function is inversely proportional it is the case of \[y \alpha \frac{ 1 }{ x }\]

OpenStudy (devonhoward15):

0.5

OpenStudy (devonhoward15):

but than its inverse

OpenStudy (ijlal):

just replace the proportionality sign and replace it with a constant \[y=k/x\]

OpenStudy (daniellelovee):

-10*5=-50 y=-50/x

OpenStudy (ijlal):

-10=k/5 so k=-50 so the equation becomes \[f(x)=-50/x\]

OpenStudy (devonhoward15):

so it is f(x)=-50/x

OpenStudy (ijlal):

yep as @Daniellelovee and me said it :)

OpenStudy (daniellelovee):

yes I think so :)